If A = {5,6,7} and B = {2,3}, {(2,6),(2,5),(3,7)} ⊆ B×A.


1
Expert's answer
2022-04-21T04:16:54-0400

Consider the Cartesian product "B\times A=\{(2,5),(2,6),(2,7),(3,5),(3,6),(3,7)\}". As we can see, the pairs "(2,6), (2,5)" and "(3,7)" belong to the set "B\times A". Thus, "\{(2,6), (2,5),(3,7)\}\sube B\times A".